OpenStudy (anonymous):

find the 15th term in the geometric sequence below. 20,10,5,2.5,...

OpenStudy (anonymous):

to find the nth of a general term of a geometric sequence, \[a _{n}=a _{1} \times r ^{n-1}\]

OpenStudy (anonymous):

\[a = 20 , r = 10/20 = 1/2 \] T15 = a . r ^ 14 = 20 x ( 1/2 ) ^ 14 = ..... !

OpenStudy (anonymous):

a(n) = the nth term of the series a(10=the 1st term of the series r=common ratio n=nth term

OpenStudy (anonymous):

@kausarsalley how can the 15th term be greater than 20 if the sequence is 20 , 10 , 2.5 ... ( notice it is decreasing ) how can you obtain 327680 ? o.O

OpenStudy (anonymous):

so in this case, since you are finding the nth term, \[a _{n}= 20 \times 0.5^{15-1}\] \[a _{n}=0.00122.....\]

OpenStudy (anonymous):

@antoni7 i made a mistake......i have made the correction..... anyway thanks for letting me know.....

OpenStudy (anonymous):

no problem.. we are all learners here